### 30 Mar 2026 v2

@@ 3.3 Comparing SupCon with Supervised Learning
Table 1 has reported the classification performance achieved by SwinT-B under the supervised learning paradigm. Figure 6 further shows the classification performance on individual L3 habitats achieved by the supervised learning by illustrating the normalised confusion matrix (CM) produced by SwinT-B. For major habitats, it is observed that the classification performance is relatively modest on grasslands, such as Neutral Grassland, Acid Grassland, and Bracken, as well as on Fen, Marsh, Swamp (FMS) belonging to the wetland. Less than 70% of test instances among these habitats are correctly predicted by SwinT-B (shown by the diagonal values in the corresponding rows of CM). This observation indicates that these habitats are very challenging to classify even using ViT (CNNs performs closely to ViT or even worse on these habitats in our experiments). We attribute this performance degradation to the visual similarity between certain habitats, as illustrated in Figure 1. For instance, Neural Grassland, which is a major L3 habitat in the CS dataset and often displays subtle visual differences to Improved Grassland, has 34% of its test instances being misclassified as Improved grassland, deteriorating the overall performance significantly. In Appendix A, we visualise the embedding clusters of different habitats. The heavily overlapped clusters further reflect their high similarity.
@@
For our best model, SwinT-B, SupCon is observed to mitigate the misclassifications between challenging habitats. Specifically, Figure 7 is a delta confusion matrix that compares the normalised confusion matrices generated by SupCon and supervised learning on major habitats. With SupCon, Improved Grassland increases its correct classifications by 6% and reduces the misclassifications on Neutral Grassland by 6%. Similarly, 5% more Neutral Grassland instances are correctly predicted and its misclassifications to Improved Grassland are reduced by 2% with SupCon. In terms of FMS, SupCon improves the proportion of correct classifications by 14% and decrease its misclassifications to Bog and Neutral Grassland by 5% and 6%, respectively. The correct predictions of Bracken are improved by 8% with SupCon, owing to a 12% decrease of misclassifications to Acid Grassland. Conversely, SupCon sacrifices classification performance of Dwarf Shrub Heath, decreasing its correct predictions by 4%, and fails to increase the correct classifications for Calcareous Grassland. This could be due to the fact that the Calcareous Grassland is a minor L3 habitat with a very limited number of samples in CS dataset for training the model, as shown in Figure 2). We provide more insights about the performance of SupCon with the embedding analysis in Appendix A.
@@ 4 Discussion
Our experiment results show that SwinT surpasses state-of-the-art CNNs, including Wide ResNet, EfficientNetV2, and ResNext, for ground-level habitat classification. SupCon training can further boosts SwinT, achieving performance comparable to human expert. Meanwhile, the study also exposes challenges in habitat classification with ground-level imagery. We provide the following insights:
@@
  • Pathways to improving ground-level habitat classification. a) The CS photos were originally taken for plot relocation, not habitat identification. Establishing clear guidelines (e.g. framing, lighting, seasonal timing) for habitat‐focused imagery should enhance the training data and reveal the true potential of AI for ground‐level habitat classification. b) Human surveyors use a rich collection of information and often quantitative data–information absent from 2D images–to determine habitats. For example, they take into consideration the geographical and climatic information, management context, soil type, species, etc. when conducting habitat assessment in the field. Additionally, they may visit the site during different seasons in order to record species growing at different times of the year. Incorporating multi-view or time-series photos, or auxiliary metadata (e.g. date, location), could narrow the remaining performance gap. c) Since habitat is often defined by characteristic species, extending our approach to recognise key flora (e.g. dominant shrubs, indicator plants) and integrating those outputs into the habitat classifier may also boost accuracy.
@@ 5 Conclusion
Our current study focuses solely on level 3 Broad Habitats within the UK (except for Bracken, which is L4). Future work should expand the model to classify fine-grained level 4 habitats with over 80 categories, where sample sizes are still small, and assess model transferability to other countries and different habitats outside of the UK. In summary, our study paves a path toward truly scalable, cost-effective habitat monitoring based on ground-level imagery by integrating state-of-the-art AI techniques with ecological expertise to support better-informed biodiversity supports and land-use decisions nationwide.

### 14 Apr 2026 v3

@@ 3.3 Comparing SupCon with Supervised Learning
Table 1 has reported the classification performance achieved by SwinT-B under the supervised learning paradigm. Figure 6 further shows the classification performance on individual L3 habitats achieved by the supervised learning by illustrating the normalised confusion matrix (CM) produced by SwinT-B. For major habitats, it is observed that the classification performance is relatively modest on grasslands, such as Neutral Grassland, Acid Grassland, and Bracken, as well as on Fen, Marsh, Swamp (FMS) belonging to the wetland. Less than 70% of test instances among these habitats are correctly predicted by SwinT-B (shown by the diagonal values in the corresponding rows of CM). This observation indicates that these habitats are very challenging to classify even using ViT (CNNs perform closely to ViT or even worse on these habitats in our experiments). We attribute this performance degradation to the visual similarity between certain habitats, as illustrated in Figure 1. For instance, Neutral Grassland, which is a major L3 habitat in the CS dataset and often displays subtle visual differences to Improved Grassland, has 34% of its test instances being misclassified as Improved Grassland, deteriorating the overall performance significantly. In Appendix A, we visualise the embedding clusters of different habitats. The heavily overlapped clusters further reflect their high similarity.
@@
For our best model, SwinT-B, SupCon is observed to mitigate the misclassifications between challenging habitats. Specifically, Figure 7 is a delta confusion matrix that compares the normalised confusion matrices generated by SupCon and supervised learning on major habitats. With SupCon, Improved Grassland increases its correct classifications by 6% and reduces the misclassifications on Neutral Grassland by 6%. Similarly, 5% more Neutral Grassland instances are correctly predicted and its misclassifications to Improved Grassland are reduced by 2% with SupCon. In terms of FMS, SupCon improves the proportion of correct classifications by 14% and decrease its misclassifications to Bog and Neutral Grassland by 5% and 6%, respectively. The correct predictions of Bracken are improved by 8% with SupCon, owing to a 12% decrease of misclassifications to Acid Grassland. Conversely, SupCon sacrifices classification performance of Dwarf Shrub Heath, decreasing its correct predictions by 4%, and fails to increase the correct classifications for Calcareous Grassland. This could be due to the fact that the Calcareous Grassland is a minor L3 habitat with a very limited number of samples in CS dataset for training the model, as shown in Figure 2. We provide more insights about the performance of SupCon with the embedding analysis in Appendix A.
@@ 4 Discussion
Our experiment results show that SwinT surpasses state-of-the-art CNNs, including Wide ResNet, EfficientNetV2, and ResNeXt, for ground-level habitat classification. SupCon training can further boost SwinT, achieving performance comparable to human experts. Meanwhile, the study also exposes challenges in habitat classification with ground-level imagery. We provide the following insights:
@@
  • Pathways to improving ground-level habitat classification. a) The CS photos were originally taken for plot relocation, not habitat identification. Establishing clear guidelines (e.g. framing, lighting, seasonal timing) for habitat‐focused imagery should enhance the training data and reveal the true potential of AI for ground‐level habitat classification. b) Human surveyors use a rich collection of information and often quantitative data—information absent from 2D images—to determine habitats. For example, they take into consideration the geographical and climatic information, management context, soil type, species, etc. when conducting habitat assessment in the field. Additionally, they may visit the site during different seasons in order to record species growing at different times of the year. Incorporating multi-view or time-series photos, or auxiliary metadata (e.g. date, location), could narrow the remaining performance gap. c) Since habitat is often defined by characteristic species, extending our approach to recognise key flora (e.g. dominant shrubs, indicator plants) and integrating those outputs into the habitat classifier may also boost accuracy.
@@ 5 Conclusion
Our current study focuses solely on level 3 Broad Habitats within the UK (except for Bracken, which is L4). Future work should expand the model to classify fine-grained level 4 habitats with over 80 categories, where sample sizes are still small, and assess model transferability to other countries and different habitats outside of the UK. In summary, our study paves a path towards truly scalable, cost-effective habitat monitoring based on ground-level imagery by integrating state-of-the-art AI techniques with ecological expertise to support better-informed biodiversity conservation and land-use decisions.